What is the molality of C2H5oh in water solution which will freeze at -10degree Celsius? 
Kf for Water=1.86

Grade:12th pass

1 Answers

Arun
5 years ago
You forgot to write Kf value of water. 
Here, you should add Kf of water = 1.86 
We know, 
Depression in freezing pont = depression Constant × molality
e.g., ∆T = Kf × m 
Here, ∆T = 10°C [ we neglected sign ]
Kf = 1.86 
Now, 10 = 1.86 × m 
m = 10/1.86 = 5.37 
Hence , molality = 5.37 m
 
Regards
Arun (askIITians forum expert)